# Second- and third-order properties of multidimensional Langevin equations

### 24 Jan 2024 v8

@@ 2.3. Probability current density and angular momentum
As noted in [15], the stochastic rotation frequencies can be experimentally measured by averaging over angular motions. We can formally prove this by working in coordinates where $\mathbf{C}=\mathds{1}$ and use Stratonovich calculus to transform to polar coordinates $(r,\phi)$. [^1] If we use Itô calculus, then $\dot{\phi}$ has a term $\propto r^{-2}$, whose expectation diverges because a 2D Gaussian has $p(r)\propto r$ as $r\to 0$, and it would appear that $\langle\dot{\phi}\rangle$ does not exist. However, using Stratonovich calculus, no such problem arises. We have:
@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is: [^2] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9].
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
In the statistical literature, the term “integrated variable” in a stochastic process refers to a variable with no stationary distribution, but whose increments, or increments of increments, etc., do possess a stationary distribution. Here, we focus on variables integrated of order one, i.e., variables whose increments possess a stationary distribution. Such variables can be used to model e.g. the position and orientation of a particle in a homogeneous, isotropic medium. The dynamics of these variables may depend on degrees of freedom represented by variables possessing a stationary distribution. [^3] Our analysis does not explicitly address ratchet models where $\langle\dot{x}\mid x\rangle$ is periodic in $x$. Such cases can be handled by introducing a complex variable $w=e^{ix}$ if the period is taken to be $2\pi$. For linear dynamics, this type of system is modeled as:
@@ 3.3. Comparison of experimental data with theoretical predictions
As an example, for a single (integrated) variable the quantity in Eq. (81) approaches as $\tau\to\infty$ the negative of the long-time diffusivity, to be compared to the short-time diffusivity.
@@ 4. Third-order properties
vanishes in the case where $\mathbf{x}$ has continuous paths. [^4] The $x^{i}$, $x^{j}$, and $x^{k}$ need not be state coordinates; they can be any functions. In such a situation, the probability density evolution satisfies a Fokker–Planck equation [3], leading to vanishing of the “cubic variation”. This gives a relation between the angular momenta, as:
@@ 6.8. Quantitative significance of third-order effects
where $d_{0}$ is the dimension of the stationary variables only and $d$ is the total dimension including integrated variables. In the above, $k$ and $k^{\prime}$ are summed over stationary variables only, whereas $i$, $j$, $i^{\prime}$, and $j^{\prime}$ are summed over all variables including integrated variables. [^6] We may however be concerned about the possibility for a non-negative diffusivity function to have a mean far less than the standard deviation of its “linear component” (scaling as $\sqrt{d}$). To this end, we may consider the function $e^{cx}$, where $c$ is a constant, for $x\sim\mathcal{N}(0,1)$. Its mean value is $\langle e^{cx}\rangle=e^{c^{2}/2}$. When performing linear regression of $e^{cx}$ with regressor $x$, the coefficient is $\langle xe^{cx}\rangle=ce^{c^{2}/2}$. We see that the ratio between the two can be made arbitrarily large.
@@ 6.9. Finite time interval “cubic variations”
and similarly for $j$ and $k$. We see that if second-order angular momenta vanish, then the “cubic variation” at finite time intervals $\tau$ vanishes, with corrections $a\mathcal{O}((a,b)^{2})\tau^{2}$ and $\mathcal{O}(\tau^{3})$. We may evaluate quantitative significance by applying the procedure for $L$ and $b$ independently.
@@ 8.3. Third-order properties
To judge quantitative significance, we first consider the quantities $\langle x^{i}x^{j}x^{k}\rangle$, $\langle x^{i}v^{j}v^{k}\rangle$, and $\langle v^{i}v^{j}v^{k}\rangle$. These are judged according to the procedure for third moments. The other quantities are angular momenta which exist only in multiple dimensions, including $\langle x^{i}x^{j}v^{k}\rangle=L(x^{i}x^{j},x^{k})/2$. For this quantity, we write $\mathbf{x}$ in terms of $\widetilde{\mathbf{x}}$ and $\widetilde{\mathbf{v}}$ and expand. We first note the magnitudes of the angular momenta:
@@
(341) $\left\langle v^{i}\frac{\mathrm{d}[v^{j},v^{k}]}{\mathrm{d}t}\right\rangle\left\langle v^{i^{\prime}}\frac{\mathrm{d}[v^{j^{\prime}},v^{k^{\prime}}]}{\mathrm{d}t}\right\rangle$ $\sim 2\mu^{2}\langle v^{i}v^{i^{\prime}}\rangle(D^{jj^{\prime}}D^{kk^{\prime}}+D^{jk^{\prime}}D^{kj^{\prime}}),$
(342) $\left\langle x^{i}\frac{\mathrm{d}[v^{j},v^{k}]}{\mathrm{d}t}\right\rangle\left\langle v^{i^{\prime}}\frac{\mathrm{d}[v^{j^{\prime}},v^{k^{\prime}}]}{\mathrm{d}t}\right\rangle$ $\sim 2\mu^{2}\langle x^{i}v^{i^{\prime}}\rangle(D^{jj^{\prime}}D^{kk^{\prime}}+D^{jk^{\prime}}D^{kj^{\prime}}).$
@@ 8.4. Dynamics on long time-scales
We now investigate the contributions of the various terms in Eq. (368) to $\langle(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ for $\tau\ll\mathcal{O}(\mu)$. [^8] If we immediately put $\tau\sim\mathcal{O}(\mu)$, then by Eqs. (161) and (162), we will introduce a cubic term in $\langle\widetilde{x}^{i}(\tau)\widetilde{x}^{j}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle-\langle\widetilde{x}^{i}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle\langle\widetilde{x}^{j}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$, which significantly complicates matters. We do this by multiplying by 1, $\widetilde{\mathbf{x}}(0)$, or $\widetilde{\mathbf{v}}(0)$, taking the expectation and integrating. First, we address the multiplication by 1. The last line of Eq. (368) contributes $\mathcal{O}(\mu^{-1}\tau)$, while the first three lines of Eq. (368) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})$. The remaining terms contribute $\mathcal{O}(\mu^{0})a\mathcal{O}(a,b)\tau$. Thus, only the last line of Eq. (368) survives. Now, we address the multiplication by $\widetilde{x}^{m}(0)$. The last line of Eq. (368) contributes $\mathcal{O}(\mu^{-1}\tau)b$ (to be calculated later), while the first three lines of Eq. (368) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})\mathcal{O}(a,b)$. The fourth line of Eq. (368) contributes:

### 02 Mar 2024 v9

@@ 2.3. Probability current density and angular momentum
As noted in [15], the stochastic rotation frequencies can be experimentally measured by averaging over angular motions. We can formally prove this by working in coordinates where $\mathbf{C}=\mathds{1}$ and use Stratonovich calculus to transform to polar coordinates $(r,\phi)$ [^1] If we use Itô calculus, then $\dot{\phi}$ has a term $\propto r^{-2}$, whose expectation diverges because a 2D Gaussian has $p(r)\propto r$ as $r\to 0$, and it would appear that $\langle\dot{\phi}\rangle$ does not exist. However, using Stratonovich calculus, no such problem arises.. We have:
@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. We explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^2] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9].:
@@ 3.1. Preliminaries: Definition, covariance functions, and detailed balance
In the statistical literature, the term “integrated variable” in a stochastic process refers to a variable with no stationary distribution, but whose increments, or increments of increments, etc., do possess a stationary distribution. Here, we focus on variables integrated of order one, i.e., variables whose increments possess a stationary distribution. Such variables can be used to model e.g. the position and orientation of a particle in a homogeneous, isotropic medium. The dynamics of these variables may depend on degrees of freedom represented by variables possessing a stationary distribution [^3] Our analysis does not explicitly address ratchet models where $\langle\dot{x}\mid x\rangle$ is periodic in $x$. Such cases can be handled by introducing a complex variable $w=e^{ix}$ if the period is taken to be $2\pi$.. For linear dynamics, this type of system is modeled as:
@@ 3.3. Comparison of experimental data with theoretical predictions
As an example, for a single (integrated) variable the quantity in Eq. (81) approaches as $\tau\to\infty$ the negative of the long-time diffusivity, to be compared to the short-time diffusivity [^4] It should be noted that the limit $\tau\to\infty$ cannot be taken for a trajectory of finite length. To illustrate this, suppose we have a trajectory of length $N+1$, and define $\Delta x_{n}:=x((n+1)\Delta t)-x(n\Delta t)$, $n=0,1,\ldots,N-1$, $\overline{\Delta x}:=\sum_{n=0}^{N-1}\Delta x_{n}/N$. We may estimate the velocity autocovariance function as $\widehat{C}(n)=\sum_{m=0}^{N-1-n}(\Delta x_{m}-\overline{\Delta x})(\Delta x_{m+n}-\overline{\Delta x})/N(\Delta t)^{2}$, $n=0,1,\ldots,N-1$ [16]. We might then estimate the long-time diffusivity by $(\Delta t)[\widehat{C}(0)/2+\sum_{n=1}^{N-1}\widehat{C}(n)]$. However, this is equal to zero. Thus, although the velocity autocovariance function is self-averaging for any given time lag, its integral is not. If we denote the upper limit of integration by $M$, then the limits $M\to\infty$, $N\to\infty$ do not commute..
@@ 4. Third-order properties
vanishes in the case where $\mathbf{x}$ has continuous paths [^5] The $x^{i}$, $x^{j}$, and $x^{k}$ need not be state coordinates; they can be any functions.. In such a situation, the probability density evolution satisfies a Fokker–Planck equation [3], leading to vanishing of the “cubic variation”. This gives a relation between the angular momenta, as:
@@ 6.8. Quantitative significance of third-order effects
where $d_{0}$ is the dimension of the stationary variables only and $d$ is the total dimension including integrated variables. In the above, $k$ and $k^{\prime}$ are summed over stationary variables only, whereas $i$, $j$, $i^{\prime}$, and $j^{\prime}$ are summed over all variables including integrated variables [^7] We may however be concerned about the possibility for a non-negative diffusivity function to have a mean far less than the standard deviation of its “linear component” (scaling as $\sqrt{d}$). To this end, we may consider the function $e^{cx}$, where $c$ is a constant, for $x\sim\mathcal{N}(0,1)$. Its mean value is $\langle e^{cx}\rangle=e^{c^{2}/2}$. When performing linear regression of $e^{cx}$ with regressor $x$, the coefficient is $\langle xe^{cx}\rangle=ce^{c^{2}/2}$. We see that the ratio between the two can be made arbitrarily large..
@@ 6.9. Finite time interval “cubic variations”
and similarly for $j$ and $k$. We see that if second-order angular momenta vanish, then the “cubic variation” at finite time intervals $\tau$ vanishes, with corrections $a\mathcal{O}((a,b)^{2})\tau^{2}$ and $\mathcal{O}(\tau^{3})$. We may evaluate quantitative significance by treating $L$ and $b$ as uncorrelated, because they have opposite behavior under time reversal.
@@ 8.3. Third-order properties
To judge quantitative significance, we first consider the quantities $\langle x^{i}x^{j}x^{k}\rangle$ and $\langle v^{i}v^{j}v^{k}\rangle$. These are judged according to the procedure for third moments. For $\langle x^{i}v^{j}v^{k}\rangle$, we may use:
(329) $\langle x^{i}v^{j}v^{k}\rangle\langle x^{i^{\prime}}v^{j^{\prime}}v^{k^{\prime}}\rangle\sim\langle x^{i}x^{i^{\prime}}\rangle\frac{\langle v^{j}v^{j^{\prime}}\rangle\langle v^{k}v^{k^{\prime}}\rangle+\langle v^{j}v^{k^{\prime}}\rangle\langle v^{k}v^{j^{\prime}}\rangle}{2}.$
The remaining quantities are angular momenta which exist only in multiple dimensions, including $\langle x^{i}x^{j}v^{k}\rangle=L(x^{i}x^{j},x^{k})/2$. For this quantity, we write $\mathbf{x}$ in terms of $\widetilde{\mathbf{x}}$ and $\widetilde{\mathbf{v}}$ and expand. We first note the magnitudes of the angular momenta:
@@
(342) $\left\langle v^{i}\frac{\mathrm{d}[v^{j},v^{k}]}{\mathrm{d}t}\right\rangle\left\langle v^{i^{\prime}}\frac{\mathrm{d}[v^{j^{\prime}},v^{k^{\prime}}]}{\mathrm{d}t}\right\rangle$ $\sim 2\mu^{2}\langle v^{i}v^{i^{\prime}}\rangle(D^{jj^{\prime}}D^{kk^{\prime}}+D^{jk^{\prime}}D^{kj^{\prime}}).$
@@ 8.4. Dynamics on long time-scales
We now investigate the contributions of the various terms in Eq. (368) to $\langle(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ for $\tau\ll\mathcal{O}(\mu)$ [^9] If we immediately put $\tau\sim\mathcal{O}(\mu)$, then by Eqs. (161) and (162), we will introduce a cubic term in $\langle\widetilde{x}^{i}(\tau)\widetilde{x}^{j}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle-\langle\widetilde{x}^{i}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle\langle\widetilde{x}^{j}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$, which significantly complicates matters.. We do this by multiplying by 1, $\widetilde{\mathbf{x}}(0)$, or $\widetilde{\mathbf{v}}(0)$, taking the expectation and integrating. First, we address the multiplication by 1. The last line of Eq. (368) contributes $\mathcal{O}(\mu^{-1}\tau)$, while the first three lines of Eq. (368) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})$. The remaining terms contribute $\mathcal{O}(\mu^{0})a\mathcal{O}(a,b)\tau$. Thus, only the last line of Eq. (368) survives. Now, we address the multiplication by $\widetilde{x}^{m}(0)$. The last line of Eq. (368) contributes $\mathcal{O}(\mu^{-1}\tau)b$ (to be calculated later), while the first three lines of Eq. (368) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})\mathcal{O}(a,b)$. The fourth line of Eq. (368) contributes: